# DBAW-PIKAN: Dynamic Balance Adaptive Weight Kolmogorov-Arnold Neural Network for Solving Partial Differential Equations

### v1

@@ Abstract
Physics-informed neural networks (PINNs) have led to significant advancements in scientific computing by integrating fundamental physical principles with advanced data-driven techniques. However, when dealing with problems characterized by multi-scale or high-frequency features, PINNs encounter persistent and severe challenges related to stiffness in gradient flow and spectral bias, which significantly limit their predictive capabilities. To address these issues, this paper proposes a Dynamic Balancing Adaptive Weighting Physics-Informed Kolmogorov-Arnold Network (DBAW-PIKAN), designed to mitigate such gradient-related failure modes and overcome the bottlenecks in function representation. The core of DBAW-PIKAN combines the Kolmogorov-Arnold network architecture, based on learnable B-splines, with an adaptive weighting strategy that incorporates a dynamic decay upper bound. Compared to baseline models, the proposed method accelerates the convergence process and improves solution accuracy by at least an order of magnitude without introducing additional computational complexity. A series of numerical benchmarks, including the Klein-Gordon, Burgers, and Helmholtz equations, demonstrate the significant advantages of DBAW-PIKAN in enhancing both accuracy and generalization performance.
@@ 1 Introduction
In the fields of natural sciences and engineering, partial differential equations (PDEs) serve as indispensable mathematical tools for describing the spatiotemporal evolution of continuous systems. By establishing quantitative relationships between unknown functions and their partial derivatives, PDEs effectively characterize the dynamic mechanisms of physical fields, providing a fundamental mathematical framework for various continuous phenomena in nature and engineering practice. PDEs exhibit exceptional descriptive capabilities across scales, ranging from macroscopic processes such as fluid dynamics, heat transfer, and electromagnetic propagation to microscopic phenomena like quantum wave function evolution. Consequently, they play a critical foundational role in diverse disciplines, including physics, engineering, materials science, biomedicine, and finance. Many classical theoretical systems are built upon PDEs, such as the Navier–Stokes equations in fluid mechanics, the Fourier heat conduction equation, the Maxwell equations in electromagnetics, and the Schrödinger equation in quantum mechanics, all of which have become core instruments for theoretical analysis and engineering applications [guo2020solving]. Since the 1990s, the primary means of solving these equations have centered on traditional numerical methods, typically represented by the finite element method (FEM) and the finite difference method (FDM).More recently, however, researchers have increasingly explored neural networks as a transformative alternative for solving such equations[lagaris1998artificial]. Although these mesh-based numerical methods have evolved into mature theoretical systems and engineering paradigms, their inherent limitations have become increasingly prominent when addressing high-dimensional problems, complex geometric boundaries, and parameter inversion tasks [weinan2021algorithms]. On one hand, the complexity of mesh generation grows exponentially with the dimensionality of the problem, leading to a sharp escalation in computational resource consumption. On the other hand, mesh adaptation and reconstruction for irregular geometries require complex preprocessing procedures, which not only increases algorithmic difficulty but may also compromise the accuracy and stability of numerical solutions due to poor mesh quality, thereby limiting their efficiency in complex application scenarios[jin2021nsfnets, ames2014numerical, bertozzi2025partial].
@@
To overcome the inherent limitations of traditional numerical methods, Physics-Informed Neural Networks (PINNs) [raissi2019physics] have been proposed as an innovative framework that integrates deep learning with physical prior knowledge. Distinct from traditional mesh-based methods, the core innovation of PINNs lies in the physics-constrained design of the loss function. Specifically, the governing PDEs, initial conditions (ICs), and boundary conditions (BCs) are embedded into the loss function as regularization terms, encoding these composite constraints as the training objective of the neural network. During the training process, the parameterized neural network $u_{\theta}$ gradually approximates the ground-truth solution of the PDE by minimizing this total loss function. The primary advantage of this approach is its mesh-free nature, which effectively alleviates the ”curse of dimensionality.” Furthermore, PINNs enable the unified modeling and solving of both forward problems (solution prediction) and inverse problems (parameter identification), significantly reducing computational complexity in complex scenarios.Since the inception of the PINN framework, it has demonstrated remarkable adaptability and potential across various multidisciplinary scientific and engineering applications. For instance, the Karniadakis group applied PINNs to the numerical simulation of complex fluid dynamics, successfully achieving efficient solutions for highly nonlinear problems such as unsteady flows and turbulence [jin2021nsfnets, raissi2019deep]. In the field of biomedical engineering, PINNs have been employed for simulating hemodynamic evolution and analyzing the mechanical properties of soft tissues, providing new technical support for clinical diagnosis and medical device design [dousse2022multi, peng2020multiscale]. Additionally, this method has been successfully applied to dynamic process simulation in geosciences [he2021physics], structural stress analysis in solid mechanics [haghighat2021physics], and multi-scale heat conduction problems [cai2021flow], fully validating its cross-disciplinary universality and engineering practicality.
Despite the widespread application of PINNs, two core bottlenecks continue to hinder their performance when solving multi-scale or highly nonlinear PDE problems. First, as pointed out by Wang et al. [wang2021understanding], the training process of PINNs suffers from ”gradient flow pathologies” arising from the underlying gradient flow dynamics. The gradient magnitudes of different components within the composite loss function can differ by several orders of magnitude. Specifically, gradients of the PDE residuals, which often involve higher-order derivatives, tend to dominate the total gradient, thereby overshadowing the gradient signals from the initial conditions (ICs) and boundary conditions (BCs). This extreme imbalance among gradient components can cause the optimization process to deviate from the optimal direction, leading the model to converge to ”spurious solutions” that satisfy the governing PDEs but severely violate physical constraints, or to undergo persistent oscillations on the loss landscape, manifesting as training instability and convergence difficulties. Therefore, dynamically coordinating the contribution weights of various loss terms during training has become a prerequisite for the successful application of PINNs.
The second core challenge stems from the inherent spectral bias [rahaman2019spectral] of Multi-Layer Perceptrons (MLPs). In the process of function approximation, MLPs exhibit a significant preference for low-frequency components, meaning they prioritize learning the smooth features of the target function while lacking sufficient capability to fit local high-frequency or multi-scale structures. In typical PDE problems such as shock wave propagation, boundary layer evolution, and stress concentration, the solution distributions often possess pronounced local high-frequency characteristics, which are difficult for standard MLPs to characterize effectively. More critically, PINNs rely on automatic differentiation (AD) to calculate PDE residuals. Higher-order differential operations further amplify errors caused by the network’s inadequate representation of high-frequency components. This leads to a significant discrepancy between the estimated and true values of the PDE residuals, preventing the optimizer from receiving effective guidance signals for training. This intrinsic mechanism fundamentally limits the convergence rate and final approximation accuracy of MLP-based PINNs.
To address the two aforementioned challenges, existing research has primarily evolved along two distinct trajectories: optimizing training dynamics and enhancing the expressive power of network architectures.
Regarding the optimization of training dynamics to mitigate gradient flow pathologies, various dynamic adaptive weighting strategies have been developed to balance the contributions of different loss terms during optimization. These approaches often draw inspiration from multi-task learning, such as employing uncertainty-based weighting [kendall2018multitask] or designing heuristic algorithms for self-adaptive weight adjustment [mcclenny2023selfadaptive, xiang2022selfadaptive].
@@
In terms of enhancing architectural expressivity, one prevalent approach involves improving input representations, such as introducing Fourier feature encoding [tancik2020fourier] to facilitate the learning of high-frequency functions. A more fundamental route lies in revolutionizing the network architecture itself. Recently, Kolmogorov-Arnold Networks (KAN), proposed by Liu et al. [liu2024kan], have emerged as a promising alternative. Unlike MLPs, KAN places learnable spline-based activation functions on the ”edges” of the network rather than using fixed activations on the ”nodes.” This design exhibits significant potential in terms of function approximation accuracy and parameter efficiency [wang2025kolmogorov], offering a viable path to overcoming the expressivity bottlenecks of MLPs.
However, these two research trajectories remain largely decoupled. The potential of an expressive architecture, such as KAN, cannot be fully realized if its training process remains unstable. Conversely, an effective weighting strategy applied to an architecture with limited expressivity, such as an MLP, will inevitably be constrained by a lower performance ceiling.
Therefore, this study proposes an integrated Dynamic Balancing Adaptive Weighting Physics-Informed Kolmogorov-Arnold Network (DBAW-PIKAN) framework, aimed at simultaneously addressing the dual challenges of architectural expressivity limitations and optimization imbalance inherent in PINNs.
@@
The core of the proposed framework lies in the adoption of KAN [liu2024kan] as a high-expressivity function approximator to overcome spectral bias, coupled with a novel DBAW (Dynamic Balancing Adaptive Weighting) strategy to stabilize the training process. The DBAW strategy constrains the self-adaptive weights by introducing a dynamically decaying upper bound, thereby effectively mitigating ”gradient flow pathologies” and preventing the weight explosion issues often encountered in adaptive weighting methods. Subsequent experimental results will demonstrate that the synergy between this architecture and the optimization strategy yields significant improvements in both solution accuracy and training stability compared to baseline models.
@@ 2 Methodology
To address the dual challenges of function approximation capability and multi-objective optimization balance encountered by Physics-Informed Neural Networks (PINNs), this study proposes an integrated DBAW-PIKAN framework. The central idea of this framework is to optimize the training process of the Kolmogorov-Arnold Network-based PIKAN architecture through the introduction of a novel dynamic adaptive weighting strategy.
This section systematically deconstructs the proposed integrated approach. Section 2.1 first reviews the standard PINN framework. Section 2.2 introduces the PIKAN framework in detail, which serves as the cornerstone of the model’s expressive power. Section 2.3 elaborates on the DBAW strategy designed to stabilize the optimization process. Finally, Section 2.4 demonstrates how these components are synergistically combined to construct the ultimate DBAW-PIKAN framework.
@@ 2.1 Physics-informed neural networks
Physics-informed neural networks aim to infer a continuous latent function $u(x,t)$ that represents the solution to a system of partial differential equations (PDEs). A general nonlinear PDE system can be expressed as:
$u_{t}+\mathcal{N}_{x}[u]=0,\quad x\in\Omega,t\in T$ (1)
where $u_{t}$ denotes the partial derivative of $u$ with respect to time $t$, and $\mathcal{N}_{x}[\cdot]$ is a linear or nonlinear differential operator acting on the spatial variable $x$ and its derivatives. To obtain a well-posed solution, the system must satisfy a set of initial conditions (ICs) and boundary conditions (BCs). The IC is defined as:
@@
and the BC is defined as:
$u(x,t)=g(x,t),\quad x\in\partial\Omega,t\in T$ (3)
In the PINN framework, the solution $u(x,t)$ is approximated by a deep neural network $u_{\theta}(x,t)$, where $\theta$ represents the trainable parameters of the network. The core of this approach lies in embedding the physical laws—including the governing PDEs, ICs, and BCs—into the network’s learning process as soft constraints. This is achieved by defining a composite loss function $\mathcal{L}(\theta)$, which typically consists of three components: the PDE residual loss ($\mathcal{L}_{r}$), the initial condition loss ($\mathcal{L}_{ic}$), and the boundary condition loss ($\mathcal{L}_{bc}$). $\mathcal{L}_{r}$ penalizes the extent to which the neural network $u_{\theta}$ violates the governing PDE within the domain, while $\mathcal{L}_{ic}$ and $\mathcal{L}_{bc}$ penalize the discrepancies between $u_{\theta}$ and the specified initial and boundary conditions, respectively. These loss terms are commonly defined using the Mean Squared Error (MSE) [Raissi2017PINNpart1]:
$\mathcal{L}_{r}=\frac{1}{N_{r}}\sum_{i=1}^{N_{r}}|r_{\theta}(x_{r}^{i},t_{r}^{i})|^{2}$ (4)
$\mathcal{L}_{bc}=\frac{1}{N_{bc}}\sum_{i=1}^{N_{bc}}|u_{\theta}(x_{bc}^{i},t_{b}^{i})-g(x_{b}^{i},t_{b}^{i})|^{2}$ (5)
$\mathcal{L}_{ic}=\frac{1}{N_{ic}}\sum_{i=1}^{N_{ic}}|u_{\theta}(x_{ic}^{i},t_{0})-h(x_{ic}^{i})|^{2}$ (6)
@@
$r_{\theta}(x,t):=\frac{\partial}{\partial t}u_{\theta}(x,t)+\mathcal{N}_{x}[u_{\theta}(x,t)]$ (7)
@@
$\mathcal{L}(\theta)=w_{r}\mathcal{L}_{r}+w_{ic}\mathcal{L}_{ic}+w_{bc}\mathcal{L}_{bc}$ (8)
The workflow of this framework is illustrated in Fig. 1. The optimization objective for the network parameters $\theta$ is to minimize this total loss function, a process typically executed via gradient-based optimization algorithms such as the Adam optimizer. In the standard PINN framework, the weights $w_{r},w_{ic},$ and $w_{bc}$ are usually treated as fixed hyperparameters, the determination of which often relies on time-consuming and empirical manual tuning. As previously noted, the significant discrepancies in magnitude and convergence rates among different loss terms constitute one of the critical challenges leading to training instability and hindering model convergence.
Figure 1: PINN Framework Diagram
@@ 2.2 Physics-Informed Kolmogorov-Arnold Network
To address the inherent limitations in the expressive power of traditional Multi-Layer Perceptrons, particularly the spectral bias problem where the network struggles to learn complex functions containing multi-scale or high-frequency information, this study introduces Kolmogorov-Arnold Networks as a high-capacity function approximator [rahaman2019spectral]. Unlike MLPs, which concentrate nonlinear transformations at nodes with fixed activation functions, KANs place learnable nonlinear activation functions on the edges of the network. The theoretical foundation of KAN is the Kolmogorov-Arnold Representation Theorem (KART)[arnold1957], which states that any multivariate continuous function $f(\mathbf{x})$ can be represented as a composition of a finite number of univariate functions and addition operations:
$f(\mathbf{x})=\sum_{q=1}^{2n+1}\Phi_{q}\left(\sum_{p=1}^{n}\phi_{q,p}(x_{p})\right)$ (9)
@@
$\psi_{j}(\mathbf{x}{\text{in}})=\sum^{i=1}{n_{\text{in}}}\phi_{j,i}(x_{i}),\quad j=1,\ldots,n_{\text{out}}$ (10)
@@
$\phi(x)=w_{b}\cdot b(x)+\sum_{k=1}^{G}c_{k}\cdot B_{k}(x)$ (11)
where $b(x)$ is a fixed basis function (e.g., SiLU) and the coefficients $c_{k}$ of the B-spline part are trainable parameters. In the context of solving PDEs, B-splines are selected due to their advantages in numerical stability, local adaptability, and derivative computation efficiency[toshniwal2020multi].Firstly, B-splines possess the local support property, meaning a single basis function is non-zero only within a finite interval. This implies that modifying a coefficient $c_{k}$ only affects the function’s shape in a local region. Since PDE solutions often feature localized sharp variations (such as shock waves or boundary layers), the locality of B-splines enables KAN to adaptively capture these phenomena with high parameter efficiency, thereby mitigating the spectral bias of MLPs.Secondly, regarding numerical stability, B-splines, as piecewise polynomials, effectively avoid the Runge phenomenon—severe oscillations near the endpoints of an interval—common in high-order global polynomial interpolation[magueresse2025energy] . In the PINN framework, unstable function approximators adversely affect the gradients computed via automatic differentiation (AD); thus, the stability of B-splines is critical for ensuring a robust optimization process. Furthermore, calculating PDE residuals requires high-order derivatives of the network output. The derivatives of B-splines have closed-form analytical expressions, allowing any order of derivative to be computed directly, accurately, and efficiently. This provides superior performance and more robust gradient estimation compared to methods relying solely on AD [beccari2022stable].
@@
$\mathbf{X}^{(l+1)}=\tanh\Bigl(\sum\phi^{(l+1)}(\mathbf{X}^{(l)})+\mathbf{W}^{(l+1)}\sigma(\mathbf{X}^{(l)})\Bigr)$ (12)
@@
$u_{\theta}(x)=(\mathcal{H}^{(L)}\circ\mathcal{H}^{(L-1)}\circ\cdots\circ\mathcal{H}^{(1)})(x)$ (13)

### v2

# Synergizing Kolmogorov-Arnold Networks with Dynamic Adaptive Weighting for High-Frequency and Multi-Scale PDE Solutions

@@ Abstract
PINNs enhance scientific computing by incorporating physical laws into neural network structures, leading to significant advancements in scientific computing. However, PINNs struggle with multi-scale and high-frequency problems due to pathological gradient flow and spectral bias, which severely limit their predictive power. By combining an enhanced network architecture with a dynamically adaptive weighting mechanism featuring upper-bound constraints, we propose the Dynamic Balancing Adaptive Weighting Physics-Informed Kolmogorov-Arnold Network (DBAW-PIKAN). The proposed method effectively mitigates gradient-related failure modes and overcomes bottlenecks in function representation. Compared to baseline models, the proposed method accelerates the convergence process and improves solution accuracy by at least an order of magnitude without introducing additional computational complexity. Numerical results on the Klein-Gordon, Burgers, and Helmholtz equations demonstrate that DBAW-PIKAN achieves superior accuracy and generalization performance.
@@ 1 Introduction
Partial differential equations (PDEs) describe the continuous evolution of physical quantities in space and time through the relationship between unknown functions and their partial derivatives[22]. The practical solution of PDEs is severely constrained by in computational efficiency and geometric adaptability with traditional mesh-based methods. Physics-Informed Neural Networks (PINNs) integrate physical constraints into the loss function, enabling solutions that possess both physical interpretability and numerical advantages. This method alleviates the challenges faced by traditional mesh-based methods in high-dimensional modeling, complex boundary handling, and parameter inversion, notably their high computational costs, difficulties in mesh adaptation, and stability problems, thereby directly addressing these bottlenecks[6]. Since the inception of the PINN framework, it has demonstrated remarkable adaptability and potential across various multidisciplinary scientific and engineering applications[12],including fluid dynamics[21], biomedical engineering, and geosciences. For instance, the Karniadakis group applied PINNs to the numerical simulation of complex fluid dynamics, successfully achieving efficient solutions for highly nonlinear problems such as unsteady flows and turbulence [11]. In the field of biomedical engineering, PINNs have been employed for simulating hemodynamic evolution and analyzing the mechanical properties of soft tissues, providing new technical support for clinical diagnosis and medical device design [7, 18]. Additionally, this method has been successfully applied to dynamic process simulation in geosciences [9], structural stress analysis in solid mechanics [8], and multi-scale heat conduction problems [3], fully validating its cross-disciplinary universality and engineering practicality.
Although widely applied, the performance of PINNs in solving multi-scale or strongly nonlinear PDE problems is constrained by two primary factors. First, Wang et al.[27] pointed out that PINNs are subject to gradient flow pathologies during training[29]. This phenomenon stems from the imbalance in gradient magnitudes across different terms of the loss function. Gradient magnitudes across components of a composite loss function can differ by several orders of magnitude, leading to gradient dominance and optimization Therefore, dynamically coordinating the contribution weights of various loss terms during training has become a prerequisite for the successful application of PINNs. Moreover, the spectral bias existing in multilayer perceptrons limits the capability of PINNs [19, 4]. This leads to a tendency to capture low-frequency macroscopic trends, while failing to resolve high-frequency details and local discontinuities in solutions to multi-scale PDEs or problems with steep gradients. To mitigate this architectural bottleneck, Kolmogorov-Arnold Networks (KANs) have emerged as a powerful alternative. Specifically, recent advancements such as Wavelet KAN (WAV-KAN) [23] integrate wavelet basis functions into the KAN framework, providing inherent multi-resolution analysis capabilities. Unlike the global activation functions in MLPs, the localized nature of wavelets allows WAV-KAN to capture high-frequency oscillations and sharp transitions naturally from an architectural perspective, effectively bypassing the spectral bias inherent in traditional neural architectures. Consequently, the computational efficiency and solution accuracy of PINNs are affected for such problems. The performance issue is further compounded by the computational mechanism of PINNs, which rely on automatic differentiation (AD) to calculate PDE residuals. The higher-order differential operations involved in AD can amplify errors stemming from the network’s inadequate representation of high-frequency components. This amplification leads to a significant discrepancy between the estimated and true residual values, which in turn compromises the optimizer and disrupts the overall training dynamics. Consequently, this intrinsic mechanism fundamentally limits the convergence rate and final approximation accuracy of MLP-based PINNs.
@@
To address the problems mentioned above, existing studies have adopted two main approaches: optimizing training dynamics and improving the representational capacity of network architectures. Regarding the optimization of training dynamics to mitigate gradient flow pathologies, various dynamic adaptive weighting strategies have been developed to balance the contributions of different loss terms during optimization. These approaches often draw inspiration from multi-task learning, such as employing uncertainty-based weighting [5] or designing heuristic algorithms for self-adaptive weight adjustment [16, 31]. In terms of enhancing architectural expressivity, one prevalent approach involves improving input representations, such as introducing Fourier feature encoding [25, 28] to facilitate the learning of high-frequency functions. However, this method does not fundamentally resolve the spectral bias of neural networks; rather, it shifts or complexifies the issue. In some cases, this can lead to training dynamics that are difficult to predict and control. Therefore, researchers are committed to pursuing fundamental solutions by innovating at the foundational level of network architecture. Recently, Kolmogorov-Arnold Networks (KAN), proposed by Liu et al. [14] represent a significant breakthrough in this direction. Unlike MLPs, KAN places learnable spline-based activation functions on the ”edges” of the network rather than using fixed activations on the ”nodes.” This design exhibits significant potential in terms of function approximation accuracy and parameter efficiency [30, 24], offering a viable path to overcoming the expressivity bottlenecks of MLPs.
@@
Adaptive weighting strategies and KAN networks primarily address the issues of pathological gradient flow and spectral bias, respectively, yet neither approach is capable of simultaneously resolving both problems. The potential of an expressive architecture, such as KAN, cannot be fully realized if its training process remains unstable. Conversely, an effective weighting strategy applied to an architecture with limited expressivity, such as an MLP, will inevitably be constrained by a lower performance ceiling. Therefore, we propose an integrated Dynamic Balancing Adaptive Weighting Physics-Informed Kolmogorov-Arnold Network (DBAW-PIKAN) framework, aimed at simultaneously addressing the dual challenges of architectural expressivity limitations and optimization imbalance inherent in PINNs. The core of the proposed framework lies in the adoption of KAN [14] as a high-expressivity function approximator to overcome spectral bias, coupled with a novel DBAW strategy to stabilize the training process. DBAW constrains adaptive weights by introducing a dynamically decaying upper bound, effectively mitigating pathological gradient flow while preventing the weight explosion risk commonly associated with traditional methods. Subsequent experimental results will demonstrate that the synergy between this architecture and the optimization strategy yields significant improvements in both solution accuracy and training stability compared to baseline models.
@@ 2 Methodology
We propose a DBAW-PIKAN framework to address the dual challenges of representational capacity and multi-objective optimization balance encountered by PINNs. A novel dynamic adaptive weighting strategy is introduced to optimize the training of the Physics-Informed Kolmogorov-Arnold Network (PIKAN) architecture, which constitutes the core of the proposed framework.
In this section, we present a systematic overview of the proposed integrated framework. Section 2.1 first reviews the standard PINNs. Section 2.2 introduces the PIKAN framework in detail, which serves as the cornerstone of the model’s expressive power. Section 2.3 elaborates on the DBAW strategy designed to stabilize the optimization process. Finally, Section 2.4 demonstrates how these components are synergistically combined to construct the ultimate DBAW-PIKAN framework.
@@ 2.1 Physics-informed neural networks
PINNs aim to infer a continuous latent function $u(x,t)$ that represents the solution to a system of partial differential equations (PDEs). A general nonlinear PDE system can be expressed as:
$u_{t}+\mathcal{N}_{x}[u]=0,\quad x\in\Omega,t\in T,$ (1)
where $u_{t}$ denotes the partial derivative of $u$ with respect to time $t$, and $\mathcal{N}_{x}[\cdot]$ is a linear or nonlinear differential operator acting on the spatial variable $x$ and its derivatives. To obtain a well-posed solution, the system must satisfy a set of initial conditions and boundary conditions. The initial conditions are defined as:
@@
and the boundary conditions are denoted as:
$u(x,t)=g(x,t),\quad x\in\partial\Omega,t\in T.$ (3)
During the training of PINNs, the solution $u(x,t)$ is approximated by a deep neural network $u_{\theta}(x,t)$, where $\theta$ represents the trainable parameters of the network. The core of this approach lies in embedding the physical laws, which including the governing PDEs, initial conditions, and boundary conditions, into the learning process of network as soft constraints. This is achieved by defining a composite loss function $\mathcal{L}(\theta)$, which typically consists of three components: the PDE residual loss ($\mathcal{L}_{r}$), the initial condition loss ($\mathcal{L}_{ic}$), and the boundary condition loss ($\mathcal{L}_{bc}$). $\mathcal{L}_{r}$ penalizes the extent to which the neural network $u_{\theta}$ violates the governing PDE within the domain, while $\mathcal{L}_{ic}$ and $\mathcal{L}_{bc}$ penalize the discrepancies between $u_{\theta}$ and the specified initial and boundary conditions, respectively. These loss terms are commonly defined using the Mean Squared Error (MSE) [10]:
$\mathcal{L}_{r}=\frac{1}{N_{r}}\sum_{i=1}^{N_{r}}|r_{\theta}(x_{r}^{i},t_{r}^{i})|^{2},$ (4)
$\mathcal{L}_{bc}=\frac{1}{N_{bc}}\sum_{i=1}^{N_{bc}}|u_{\theta}(x_{bc}^{i},t_{b}^{i})-g(x_{b}^{i},t_{b}^{i})|^{2},$ (5)
$\mathcal{L}_{ic}=\frac{1}{N_{ic}}\sum_{i=1}^{N_{ic}}|u_{\theta}(x_{ic}^{i},t_{0})-h(x_{ic}^{i})|^{2},$ (6)
@@
$r_{\theta}(x,t):=\frac{\partial}{\partial t}u_{\theta}(x,t)+\mathcal{N}_{x}[u_{\theta}(x,t)].$ (7)
@@
$\mathcal{L}(\theta)=w_{r}\mathcal{L}_{r}+w_{ic}\mathcal{L}_{ic}+w_{bc}\mathcal{L}_{bc}.$ (8)
The workflow of this framework is illustrated in Fig. 1. The optimization objective for the network parameters $\theta$ is to minimize this total loss function, a process typically executed via gradient-based optimization algorithms such as the Adam optimizer. In the standard PINNs framework, the weights $w_{r},w_{ic},$ and $w_{bc}$ are usually treated as fixed hyperparameters, the determination of which often relies on time-consuming and empirical manual tuning. As previously noted, the significant discrepancies in magnitude and convergence rates among different loss terms constitute one of the critical challenges leading to training instability and hindering model convergence.
Figure 1: PINNs Framework Diagram
@@ 2.2 Physics-Informed Kolmogorov-Arnold Network
To address the inherent limitations in the expressive power of traditional Multi-Layer Perceptrons, particularly the spectral bias problem where the network struggles to learn complex functions containing multi-scale or high-frequency information, we introduce Kolmogorov-Arnold Networks as a high-capacity function approximator [19]. Unlike MLPs, which concentrate nonlinear transformations at nodes with fixed activation functions, KANs place learnable nonlinear activation functions on the edges of the network. The theoretical foundation of KAN is the Kolmogorov-Arnold Representation Theorem (KART)[1], which states that any multivariate continuous function $f(\mathbf{x})$ can be represented as a composition of a finite number of univariate functions and addition operations:
$f(\mathbf{x})=\sum_{q=1}^{2n+1}\Phi_{q}\left(\sum_{p=1}^{n}\phi_{q,p}(x_{p})\right).$ (9)
@@
$\psi_{j}(\mathbf{x}_{\text{in}})=\sum^{i=1}{n_{\text{in}}}\phi_{j,i}(x_{i}),\quad j=1,\ldots,n_{\text{out}},$ (10)
@@
$\phi(x)=w_{b}\cdot b(x)+\sum_{k=1}^{G}c_{k}\cdot B_{k}(x),$ (11)
where $b(x)$ is a fixed basis function (e.g., SiLU) and the coefficients $c_{k}$ of the B-spline part are trainable parameters. In the context of solving PDEs, B-splines are selected due to their advantages in numerical stability, local adaptability, and derivative computation efficiency[26]. Firstly, B-splines possess the local support property, meaning a single basis function is non-zero only within a finite interval. This implies that modifying a coefficient $c_{k}$ only affects the function’s shape in a local region. Since PDE solutions often feature localized sharp variations (such as shock waves or boundary layers), the locality of B-splines enables KAN to adaptively capture these phenomena with high parameter efficiency, thereby mitigating the spectral bias of MLPs. Additionally, regarding numerical stability, B-splines, as piecewise polynomials, effectively avoid the Runge phenomenon, which refers to severe oscillations near the endpoints of an interval and is common in high-order global polynomial interpolation.[15] . In the PINNs framework, unstable function approximators adversely affect the gradients computed via automatic differentiation (AD); thus, the stability of B-splines is critical for ensuring a robust optimization process. Furthermore, calculating PDE residuals requires high-order derivatives of the network output. The derivatives of B-splines have closed-form analytical expressions, allowing any order of derivative to be computed directly, accurately, and efficiently. This provides superior performance and more robust gradient estimation compared to methods relying solely on AD [2].
@@
$\mathbf{X}^{(l+1)}=\tanh\Bigl(\sum\phi^{(l+1)}(\mathbf{X}^{(l)})+\mathbf{W}^{(l+1)}\sigma(\mathbf{X}^{(l)})\Bigr),$ (12)
@@
$u_{\theta}(x)=(\mathcal{H}^{(L)}\circ\mathcal{H}^{(L-1)}\circ\cdots\circ\mathcal{H}^{(1)})(x).$ (13)